# A Context for Manifold Calculus

### v2

@@ 3.1. Statement of the Main Result
Let $X$ be a topological space, and let $k\geq 0$. A subset $\mathcal{U}\subset\operatorname{Open}\left(X\right)$ of $X$ is called a Weiss $k$-cover if for each subset $S\subset X$ of cardinality $\leq k$, there is an element of $\mathcal{U}$ which contains $S$. The Weiss $k$-topology on $X$ is the Grothendieck topology on $N\left(\operatorname{Open}\left(X\right)\right)$ whose covering sieves of each open set $U\in\operatorname{Open}\left(X\right)$ are precisely the Weiss $k$-covers $\mathcal{U}$ of $U$ that are also sieves on $U$. If $\mathcal{C}$ is an $\infty$-category, a functor $F:N\left(\operatorname{Open}\left(X\right)\right)\to\mathcal{C}$ is called a Weiss $k$-cosheaf if it is a cosheaf with respect to the Weiss $k$-topology. The Weiss topology on $X$ is the intersection of the Weiss $k$-topologies as $k$ ranges over all nonnegative integers. A Weiss cosheaf is a cosheaf with respect to the Weiss topology.
@@
Let $X$ be a topological space and $\mathcal{C}$ an $\infty$-category.
@@
    Any nonempty collection of open sets of $X$ is a Weiss 00-cover of $X$. Consequently, a functor $F:N\left(\operatorname{Open}\left(X\right)\right)\to\mathcal{C}$ is a Weiss 00-cosheaf if and only if it is essentially constant, i.e., it factors through a contractible Kan complex. Since $N\left(\operatorname{Open}\left(X\right)\right)$ is already weakly contractible, this is equivalent to the requirement that $F$ carries all morphisms to equivalences.
@@
    A functor $F:N\left(\operatorname{Open}\left(X\right)\right)\to\mathcal{C}$ is a Weiss $1$-cosheaf if and only if for each nonempty sieve $\mathcal{U}\subset\operatorname{Open}\left(X\right)$, the map
@@
    is an equivalence. Thus a Weiss $1$-cosheaf is a cosheaf on $X$ (that is, a cosheaf with respect to the standard Grothendieck topology on $\operatorname{Open}\left(X\right)$) if and only if $F\left(\emptyset\right)$ is an initial object.
@@
Let $k\geq 0$, let $X$ be a topological space, and let $\chi:\mathcal{I}\to N\left(\operatorname{Open}\left(X\right)\right)$ be a functor of small $\infty$-categories. For each finite set $S\subset X$, let $\mathcal{I}_{S}\subset\mathcal{I}$ denote the full subcategory spanned by the objects $I\in\mathcal{I}$ such that $S\subset\chi\left(I\right)$.
@@
We now arrive at the identification result of polynomial functors with cosheaves and Kan extensions, which we prove in Subsection 3.3.
@@
Let $k\geq 0$, let $M$ be a manifold, let $\mathcal{C}$ be an $\infty$-category with small colimits, and let $F:N\left(\operatorname{Open}\left(M\right)\right)\to\mathcal{C}$ be an isotopy functor. The following conditions are equivalent:
@@
    The functor $F$ is a left Kan extension of $F|N\left(\operatorname{Disj}^{\leq k}\left(M\right)\right)$.
@@
    Let $\mathcal{I}$ be a small $\infty$-category, let $U\subset M$ be an open set, and let $\chi:\mathcal{I}\to N\left(\operatorname{Open}\left(U\right)\right)$ be a functor satisfying the Weiss $k$-condition. Then the map
@@
Let $M$ be a manifold, let $\mathcal{C}$ be an $\infty$-category with small colimits, and let $F:N\left(\operatorname{Open}\left(M\right)\right)\to\mathcal{C}$ be an isotopy functor. The following conditions are equivalent:
@@
    The functor $F$ is a left Kan extension of $F|N\left(\operatorname{Disj}\left(M\right)\right)$.
@@
    Let $\mathcal{I}$ be a small $\infty$-category, let $U\subset M$ be an open set, and let $\chi:\mathcal{I}\to N\left(\operatorname{Open}\left(U\right)\right)$ be a functor which satisfies the Weiss condition. Then the map
@@
Assuming Lemma 3.5 and Theorem 3.10, we can prove Theorem 3.4 as follows:
@@
$L:\operatorname{Fun}_{\mathrm{istp}}\left(N\left(\operatorname{Disj}^{\leq k}\left(M\right)\right),\mathcal{C}\right)\to\operatorname{Fun}_{\mathrm{istp}}\left(N\left(\operatorname{Open}\left(M\right)\right),\mathcal{C}\right),$
@@
3.2. Proof of Lemma 3.5
In this subsection, we prove Lemma 3.5.
Notation 3.12.
Let $n,k\geq 0$, let $M$ be an $n$-manifold, and let $\mathcal{C}$ be an $\infty$-category with small colimits. We consider the following $\infty$-categories.
  1. (1)
    The full subcategory $\operatorname{Fun}^{\leq k}\left(N\left(\operatorname{Open}\left(M\right)\right),\mathcal{C}\right)\subset\operatorname{Fun}\left(N\left(\operatorname{Open}\left(M\right)\right),\mathcal{C}\right)$ spanned by the functors that are left Kan extensions of their restrictions to $N\left(\operatorname{Disj}^{\leq k}\left(M\right)\right)$.
  2. (2)
    The full subcategory $\operatorname{Fun}^{\infty}\left(N\left(\operatorname{Open}\left(M\right)\right),\mathcal{C}\right)\subset\operatorname{Fun}\left(N\left(\operatorname{Open}\left(M\right)\right),\mathcal{C}\right)$ spanned by the functors that are left Kan extensions of their restrictions to $N\left(\operatorname{Disj}\left(M\right)\right)$.
  3. (3)
    We let $\operatorname{Fun}_{\mathrm{istp}}^{\infty}\left(N\left(\operatorname{Open}\left(M\right)\right),\mathcal{C}\right)$ denote the intersection of $\operatorname{Fun}_{\mathrm{istp}}\left(N\left(\operatorname{Open}\left(M\right)\right),\mathcal{C}\right)$ and $\operatorname{Fun}^{\infty}\left(N\left(\operatorname{Open}\left(M\right)\right),\mathcal{C}\right)$, and define $\operatorname{Fun}_{\mathrm{istp}}^{\leq k}\left(N\left(\operatorname{Open}\left(M\right)\right),\mathcal{C}\right)$ similarly.
  4. (4)
    The full subcategory $\operatorname{Fun}^{\leq k}\left(\mathcal{M}\mathsf{fld}_{n/M},\mathcal{C}\right)\subset\operatorname{Fun}\left(\mathcal{M}\mathsf{fld}_{n/M},\mathcal{C}\right)$ spanned by the functors that are left Kan extensions of their restrictions to $\mathcal{D}\mathsf{isk}_{n/M}^{\leq k}$.
  5. (5)
    The full subcategory $\operatorname{Fun}^{\infty}\left(\mathcal{M}\mathsf{fld}_{n/M},\mathcal{C}\right)\subset\operatorname{Fun}\left(\mathcal{M}\mathsf{fld}_{n/M},\mathcal{C}\right)$ spanned by the functors that are left Kan extensions of their restrictions to $\mathcal{D}\mathsf{isk}_{n/M}$.
Lemma 3.13.
Let $n,k\geq 0$, let $M$ be an $n$-manifold, and let $\mathcal{C}$ be an $\infty$-category with small colimits. Then:
  1. (1)
    The functor $\operatorname{Fun}^{\infty}\left(\mathcal{M}\mathsf{fld}_{n/M},\mathcal{C}\right)\to\operatorname{Fun}_{\mathrm{istp}}^{\infty}\left(N\left(\operatorname{Open}\left(M\right)\right),\mathcal{C}\right)$ is well-defined and is a categorical equivalence.
  2. (2)
    The functor $\operatorname{Fun}^{\leq k}\left(\mathcal{M}\mathsf{fld}_{n/M},\mathcal{C}\right)\to\operatorname{Fun}_{\mathrm{istp}}^{\leq k}\left(N\left(\operatorname{Open}\left(M\right)\right),\mathcal{C}\right)$ is well-defined and is a categorical equivalence.
  3. (3)
    A functor $F\in\operatorname{Fun}^{\infty}\left(N\left(\operatorname{Open}\left(M\right)\right),\mathcal{C}\right)$ is isotopy invariant if and only if $F|N\left(\operatorname{Disj}\left(M\right)\right)$ is isotopy invariant.
  4. (4)
    A functor $F\in\operatorname{Fun}^{\leq k}\left(N\left(\operatorname{Open}\left(M\right)\right),\mathcal{C}\right)$ is isotopy invariant if and only if $F|N\left(\operatorname{Disj}^{\leq k}\left(M\right)\right)$ is isotopy invariant.
Proof.
We will prove assertions (1) and (3); assertions (2) and (4) can be proved similarly.
First we prove the “well-defined” part of (1). Let $F:\mathcal{M}\mathsf{fld}_{n/M}\to\mathcal{C}$ be a functor which is a left Kan extension of $F|\mathcal{D}\mathsf{isk}_{n/M}$. We wish to show that the functor $F|N\left(\operatorname{Open}\left(M\right)\right)$ is a left Kan extension of $N\left(\operatorname{Disj}\left(M\right)\right)$ and that it is isotopy invariant. The latter assertion is obvious. For the former, let $U\subset M$ be an open set. We must show that the composite
[…]
is a colimit diagram. Let $\iota:U\to M$ denote the inclusion. By hypothesis, the composite
[…]
is a colimit diagram. It will therefore suffice to show that the functor
$\theta:N\left(\operatorname{Disj}\left(M\right)\right)_{/U}\to\left(\mathcal{D}\mathsf{isk}_{n/M}\right)_{/\iota}$
is final. Since the functor $\left(\mathcal{D}\mathsf{isk}_{n/M}\right)_{/\iota}\to\mathcal{D}\mathsf{isk}_{n/U}$ is a trivial fibration, it suffices to show that the functor $N\left(\operatorname{Disj}\left(M\right)\right)_{/U}\cong N\left(\operatorname{Disj}\left(U\right)\right)\to\mathcal{D}\mathsf{isk}_{n/U}$ is final. This follows from Theorem 2.24, since localization functors are final [Lur24, Tag 02N9].
We now prove (3). We must show that for any isotopy invariant functor $F:N\left(\operatorname{Disj}\left(M\right)\right)\to\mathcal{C}$, its left Kan extension is also isotopy invariant. By Theorem 2.24, we may assume that $F$ extends to a functor $G:\mathcal{D}\mathsf{isk}_{n/M}\to\mathcal{C}$. Let $G^{\prime}:\mathcal{M}\mathsf{fld}_{n/M}\to\mathcal{C}$ denote a left Kan extension of $G$. By the result in the previous paragraph, $G^{\prime}|N\left(\operatorname{Open}\left(M\right)\right)$ is a left Kan extension of $F=G^{\prime}|N\left(\operatorname{Disj}\left(M\right)\right)$. Since $G^{\prime}|N\left(\operatorname{Open}\left(M\right)\right)$ is isotopy invariant, we are done.
Finally, we complete the proof of (1). We must show that the functor
$f:\operatorname{Fun}^{\infty}\left(\mathcal{M}\mathsf{fld}_{n/M},\mathcal{C}\right)\to\operatorname{Fun}_{\mathrm{istp}}^{\infty}\left(N\left(\operatorname{Open}\left(M\right)\right),\mathcal{C}\right)$
is a categorical equivalence. Consider the commutative diagram
${{\operatorname{Fun}^{\infty}(\mathcal{M}\mathsf{fld}_{n/M},\mathcal{C})}}$${{\operatorname{Fun}^{\infty}_{\mathrm{istp}}(N(\operatorname{Open}(M)),\mathcal{C})}}$${{\operatorname{Fun}(\mathcal{D}\mathsf{isk}_{n/M},\mathcal{C})}}$${{\operatorname{Fun}_{\mathrm{istp}}(N(\operatorname{Disj}(M)),\mathcal{C}).}}$$\scriptstyle{g}$$\scriptstyle{f}$$\scriptstyle{g^{\prime}}$$\scriptstyle{f^{\prime}}$
The functors $g$ and $g^{\prime}$ are trivial fibrations by [Lur09, Proposition 4.3.2.15], and the functor $f^{\prime}$ is a categorical equivalence by Theorem 2.24. Hence $f$ is also a categorical equivalence. The proof is now complete. ∎
Proof of Lemma 3.5.
This follows from Lemma 3.13. ∎
3.3. Proofs of Theorems 3.10 and 3.11
We now turn to the proofs of Theorems 3.10 and 3.11. We need a few lemmas.
@@
Let $n,k\geq 0$, let $\mathcal{I}$ be a (small) category and let $\overline{f}:\mathcal{I}^{\triangleright}\to\mathcal{M}\mathsf{fld}_{n}^{\Delta}$ be a functor. Set $\overline{f}\left(I\right)=U_{I}$ for $I\in\mathcal{I}$ and $\overline{f}\left(\infty\right)=M$. For each subset $S\subset M$, let $\mathcal{I}_{S}\subset\mathcal{I}$ denote the full subcategory spanned by the objects $I\in\mathcal{I}$ such that $U_{I}$ contains $S$.
@@
    Suppose that $U_{I}\in\mathcal{D}\mathsf{isk}_{n}$ for every $I\in\mathcal{I}$. If $\mathcal{I}_{S}$ is weakly contractible for every finite set $S\subset M$, then the functor $N\left(\mathcal{I}\right)\to\mathcal{D}\mathsf{isk}_{n}{}_{/M}$ is final.
@@
    Suppose that $U_{I}\in\mathcal{D}\mathsf{isk}_{n}^{\leq k}$ for every $I\in\mathcal{I}$. If $\mathcal{I}_{S}$ is weakly contractible for every finite set $S\subset M$ of cardinality $\leq k$, then the functor $N\left(\mathcal{I}\right)\to\mathcal{D}\mathsf{isk}_{n/M}^{\leq k}$ is final.
@@
We prove part (1); the proof of part (2) is similar. By Proposition 2.17 and [Lur09, Theorem 4.2.4.1], it will suffice to show that, for each object $V\in\mathcal{D}\mathsf{isk}_{n}$, the map
@@
is a weak homotopy equivalence. If $V=\emptyset$, the claim is obvious because $\mathcal{I}=\mathcal{I}_{\emptyset}$ is weakly contractible by hypothesis. So suppose that $V$ is nonempty. Let $p$ denote the cardinality of the set of components of $V$ and fix a homeomorphism $V\cong\mathbb{R}^{n}\times\{1,\dots,p\}$. By Proposition 2.9, the evaluation at the origin of $\mathbb{R}^{n}$ determines a homotopy cartesian square
@@
[…]
@@
Let $n\geq 1$, $k\geq 0$, and let $-1<c_{0}<\cdots<c_{k}<1$ and $r>0$ be real numbers. Set
@@
Figure 3.1. Picture of Lemma 3.15.
@@
    The function $\phi$ agrees with $-x^{2}$ outside $[-1,1]$
@@
    We have $-1<\phi\left(a_{1}\right)<\cdots<\phi\left(a_{k}\right)$.
@@
We then define $F:\mathbb{R}^{n}\to\mathbb{R}$ by $F\left(x\right)=\phi\left(x_{1}\right)-\sum_{1<i\leq n}x_{i}^{2}$. By construction, the critical points of $F$ in $F^{-1}\left((-\infty,r^{2}]\right)$ are the points $\{\left(a_{i},0,\dots,0\right)\}_{1\leq i\leq k}$, and all of them are nondegenerate and have index $n-1$. Therefore, the Morse theory (see, e.g., [Kos93, Proposition VII.2.2]) shows that the set $\mathbb{R}^{n}\setminus\operatorname{Int}\left(C_{0}\cup\cdots\cup C_{k}\right)=F^{-1}\left((-\infty,-r^{2}]\right)$ is obtained from $F^{-1}\left((-\infty,-1]\right)=\{x\in\mathbb{R}^{n}\mid\left\|x\right\|\geq 1\}$ by attaching $\left(n-1\right)$-handles $k$ times. ∎
@@
Let $n,k\geq 0$ and let $M$ be an $n$-manifold. Let $\mathcal{U}$ be a set of open sets of $M$ which satisfies the following conditions:
@@
    be a handle decomposition of $N$ of type $\left(a_{0},\dots,a_{n}\right)$, where we wrote $a=\sum_{i=0}^{n}a_{i}$. Since attaching a 00-handle is equivalent to adding a disjoint copy of $D^{n}$, by rearranging the order of the handle attachment, we may assume that $N_{a}$ is obtained from $N_{a-1}$ by attaching a handle of positive index, say $\lambda$. Let $e:D^{\lambda}\times D^{n-\lambda}\to N_{a}$ denote the corresponding embedding. Choose disjoint closed disks $C_{0},\dots,C_{k}\subset\operatorname{Int}D^{\lambda}$ and set $A_{j}=C_{j}\times D^{n-\lambda}$. If $S\subset\{0,\dots,k\}$ is a nonempty subset, then we have
@@
    We finish off the proof by considering the case where $n=4$. Let $U\subset M$ be a nonempty finite set. For each path component $V\subset U$, choose $\left(k+1\right)$ points $p_{0}^{V},\dots,p_{k}^{V}\in V$, and set $A_{j}=\{p_{j}^{V}\mid V\in\pi_{0}\left(U\right)\}$ for $0\leq j\leq k$. Then $A_{0},\dots,A_{k}$ are disjoint closed subsets of $U$, and moreover for each nonempty finite set $S\subset\{0,\dots,k\}$, the components of the open set $U\setminus\bigcup_{i\in S}A_{i}$ are noncompact. Since connected, non-compact $4$-manifolds are smoothable [FQ90, 8.2], we deduce from Step 2 that $\bigcap_{i\in S}U\setminus A_{i}$ belongs to $\mathcal{U}$. Condition (3) now shows that $U$ belongs to $\mathcal{U}$, and the proof is complete.
@@
Our proof of Lemma 3.16 is a modification of Weiss’s proof of [Wei99, Theorem 5.1], where he essentially established the lemma in the case where $M$ is smoothable.
We now arrive at the proof of Theorems 3.10 and 3.11.
@@
Clearly (2$\implies$(3$\implies$(4). It will therefore suffice to show that (1$\implies$(2) and (4$\implies$(1).
First we show that (1$\implies$(2). Suppose that $F$ is a left Kan extension of $F|N\left(\operatorname{Disj}^{\leq k}\left(M\right)\right)$. Let $\mathcal{I}^{\prime}$ be a small $\infty$-category, let $U\subset M$ be an open set, and let $\chi:\mathcal{I}^{\prime}\to N\left(\operatorname{Open}\left(U\right)\right)$ be a functor satisfying the Weiss $k$-condition. For each $I\in\mathcal{I}^{\prime}$, set $U_{I}=\chi\left(I\right)$. We must show that the map
$\operatorname{colim}_{I\in\mathcal{I}^{\prime}}F\left(U_{I}\right)\to F\left(U\right)$

### v3

@@ 1. Polynomial Functors
Let $M$ be a manifold, and let $k\geq 0$. A subset $\mathcal{U}\subset\operatorname{Open}\left(M\right)$ of $M$ is called a Weiss $k$-cover if for each subset $S\subset M$ of cardinality $\leq k$, there is an element of $\mathcal{U}$ which contains $S$. The Weiss $k$-topology on $M$ is the Grothendieck topology on $\operatorname{Open}\left(M\right)$ whose covering sieves of each open set $U\in\operatorname{Open}\left(M\right)$ are precisely the Weiss $k$-covers $\mathcal{U}$ of $U$ that are also sieves on $U$. If $\mathcal{C}$ is an $\infty$-category, a functor $F:\operatorname{Open}\left(M\right)\to\mathcal{C}$ is called a Weiss $k$-cosheaf if it is a cosheaf with respect to the Weiss $k$-topology. The Weiss topology on $M$ is the intersection of the Weiss $k$-topologies as $k$ ranges over all nonnegative integers. A Weiss cosheaf is a cosheaf with respect to the Weiss topology.
@@
Let $M$ be a manifold, and let $\mathcal{C}$ be an $\infty$-category.
@@
    Any nonempty collection of open sets of $M$ is a Weiss 0-cover of $X$. Consequently, a functor $F:\operatorname{Open}\left(M\right)\to\mathcal{C}$ is a Weiss 0-cosheaf if and only if it is essentially constant, i.e., it factors through a contractible Kan complex. Since $\operatorname{Open}\left(M\right)$ is already weakly contractible, this is equivalent to the requirement that $F$ carries all morphisms to equivalences.
@@
    A functor $F:\operatorname{Open}\left(M\right)\to\mathcal{C}$ is a Weiss $1$-cosheaf if and only if for each nonempty sieve $\mathcal{U}\subset\operatorname{Open}\left(M\right)$, the map
@@
    is an equivalence. Thus a Weiss $1$-cosheaf is a cosheaf on $X$ (that is, a cosheaf with respect to the standard Grothendieck topology on $\operatorname{Open}\left(M\right)$) if and only if $F\left(\emptyset\right)$ is an initial object.
@@
Let $k\geq 0$, let $M$ be a manifold, and let $\chi:\mathcal{I}\to\operatorname{Open}\left(M\right)$ be a functor of small $\infty$-categories. For each finite set $S\subset M$, let $\mathcal{I}_{S}\subset\mathcal{I}$ denote the full subcategory spanned by the objects $I\in\mathcal{I}$ such that $S\subset\chi\left(I\right)$.
@@
We now arrive at the identification result of polynomial functors with cosheaves and Kan extensions, which we prove at the end of this section.
@@
Let $k\geq 0$, let $M$ be a manifold, let $\mathcal{C}$ be an $\infty$-category with small colimits, and let $F:\operatorname{Open}\left(M\right)\to\mathcal{C}$ be an isotopy functor. The following conditions are equivalent:
@@
    The functor $F$ is a left Kan extension of $F|\operatorname{Disj}^{\leq k}\left(M\right)$.
@@
    Let $\mathcal{I}$ be a small $\infty$-category, let $U\subset M$ be an open set, and let $\chi:\mathcal{I}\to\operatorname{Open}\left(U\right)$ be a functor satisfying the Weiss $k$-condition. Then the map
@@
Let $M$ be a manifold, let $\mathcal{C}$ be an $\infty$-category with small colimits, and let $F:\operatorname{Open}\left(M\right)\to\mathcal{C}$ be an isotopy functor. The following conditions are equivalent:
@@
    The functor $F$ is a left Kan extension of $F|\operatorname{Disj}\left(M\right)$.
@@
    Let $\mathcal{I}$ be a small $\infty$-category, let $U\subset M$ be an open set, and let $\chi:\mathcal{I}\to\operatorname{Open}\left(U\right)$ be a functor which satisfies the Weiss condition. Then the map
@@
Using Lemma 1.4 and Theorem 1.9, we can prove Theorem 1.3 as follows:
@@
$L:\operatorname{Fun}_{\mathrm{istp}}\left(\operatorname{Disj}^{\leq k}\left(M\right),\mathcal{C}\right)\to\operatorname{Fun}_{\mathrm{istp}}\left(\operatorname{Open}\left(M\right),\mathcal{C}\right),$
@@
We now turn to the proofs of Theorems 1.9 and 1.10. We need a few lemmas.
@@
Let $n,k\geq 0$, let $\mathcal{I}$ be a (small) category and let $\overline{f}:\mathcal{I}^{\triangleright}\to\mathcal{M}\mathsf{fld}_{n}$ be a functor. Set $\overline{f}\left(I\right)=U_{I}$ for $I\in\mathcal{I}$ and $\overline{f}\left(\infty\right)=M$. For each subset $S\subset M$, let $\mathcal{I}_{S}\subset\mathcal{I}$ denote the full subcategory spanned by the objects $I\in\mathcal{I}$ such that $U_{I}$ contains $S$.
@@
    Suppose that $U_{I}\in\mathcal{D}\mathsf{isk}_{n}$ for every $I\in\mathcal{I}$. If $\mathcal{I}_{S}$ is weakly contractible for every finite set $S\subset M$, then the functor $\mathcal{I}\to\mathcal{D}\mathsf{isk}_{n}{}_{/M}$ is final.
@@
    Suppose that $U_{I}\in\mathcal{D}\mathsf{isk}_{n}^{\leq k}$ for every $I\in\mathcal{I}$. If $\mathcal{I}_{S}$ is weakly contractible for every finite set $S\subset M$ of cardinality $\leq k$, then the functor $\mathcal{I}\to\mathcal{D}\mathsf{isk}_{n/M}^{\leq k}$ is final.
@@
We prove part (1); the proof of part (2) is similar. By Proposition B.3 and [Lur09, Theorem 4.2.4.1], it will suffice to show that, for each object $V\in\mathcal{D}\mathsf{isk}_{n}$, the map
@@
is a weak homotopy equivalence of simplicial sets. If $V=\emptyset$, the claim is obvious because $\mathcal{I}=\mathcal{I}_{\emptyset}$ is weakly contractible by hypothesis. So suppose that $V$ is nonempty. Let $p$ denote the cardinality of the set of components of $V$, and fix a homeomorphism $V\cong\mathbb{R}^{n}\times\{1,\dots,p\}$. By Proposition C.9, the evaluation at the origin of $\mathbb{R}^{n}$ determines a homotopy cartesian square
@@
Our next lemma is completely elementary in the topological case and the PL case, but we include a proof using Morse theory because we need it in the smooth version of Lemma 1.13.
@@
Let $n\geq 1$, $k\geq 0$, and let $1<c_{0}<\cdots<c_{k}<1$ and $r>0$ be real numbers. Set
@@
Figure $1.$ Picture of Lemma 1.12.
@@
    The function $\phi$ agrees with $x^{2}$ outside [1,1][-1,1]
@@
    We have $1<\phi\left(a_{1}\right)<\cdots<\phi\left(a_{k}\right)$.
@@
We then define $F:\mathbb{R}^{n}\to\mathbb{R}$ by $F\left(x\right)=\phi\left(x_{1}\right)-\sum_{1<i\leq n}x_{i}^{2}$. By construction, the critical points of $F$ in $F^{-1}\left((-\infty,r^{2}]\right)$ are the points $\{\left(a_{i},0,\dots,0\right)\}_{1\leq i\leq k}$, and all of them are nondegenerate and have index n1n-1. Therefore, Morse theory (see, e.g., [Kos93, Proposition VII.2.2]) shows that the set $\mathbb{R}^{n}\setminus\operatorname{Int}\left(C_{0}\cup\cdots\cup C_{k}\right)=F^{-1}\left((-\infty,-r^{2}]\right)$ is obtained from $F^{-1}\left((-\infty,-1]\right)=\{x\in\mathbb{R}^{n}\mid\left\|x\right\|\geq 1\}$ by attaching $\left(n-1\right)$-handles $k$ times. ∎
[…]
@@
Let $n,k\geq 0$, and let $M$ be an $n$-manifold. Let $\mathcal{U}$ be a set of open sets of $M$ which satisfies the following conditions:
@@
    be a handle decomposition of $N$ of type $\left(a_{0},\dots,a_{n}\right)$, where we wrote $a=\sum_{i=0}^{n}a_{i}$. Since attaching a 0-handle is equivalent to adding a disjoint copy of $D^{n}$, by rearranging the order of the handle attachment, we may assume that $N_{a}$ is obtained from $N_{a-1}$ by attaching a handle of positive index, say $\lambda$. Let $e:D^{\lambda}\times D^{n-\lambda}\to N_{a}$ denote the corresponding embedding. Choose disjoint closed disks $C_{0},\dots,C_{k}\subset\operatorname{Int}D^{\lambda}$ and set $A_{j}=C_{j}\times D^{n-\lambda}$. If $S\subset\{0,\dots,k\}$ is a nonempty subset, then we have
@@
    We finish off the proof by considering the case where $n=4$. Let $U\subset M$ be a nonempty finite set. For each path component $V\subset U$, choose $\left(k+1\right)$ points $p_{0}^{V},\dots,p_{k}^{V}\in V$, and set $A_{j}=\{p_{j}^{V}\mid V\in\pi_{0}\left(U\right)\}$ for $0\leq j\leq k$. Then $A_{0},\dots,A_{k}$ are disjoint closed subsets of $U$, and moreover for each nonempty finite set $S\subset\{0,\dots,k\}$, the components of the open set $U\setminus\bigcup_{i\in S}A_{i}$ are noncompact. Since connected, non-compact $4$-manifolds are smoothable [FQ90, 8.2], we deduce from Step 2 that $\bigcap_{i\in S}U\setminus A_{i}$ belongs to $\mathcal{U}$. Condition (3) now shows that $U$ belongs to $\mathcal{U}$, and the proof is complete.∎
@@
Our proof of Lemma 1.13 is a modification of Weiss’s proof of [Wei99, Theorem 5.1], where he essentially established the lemma in the case where $M$ is smoothable. The topological case is a bit harder because we do not have a control of the order of the indices of handles.
Remark 1.15.
We have stated Lemma 1.13 to highlight its relation to exhaustivity and excisivity of functors. However, a closer inspection of the proof actually shows that we can replace condition (2) as follows:
  • (2)
    Let $U_{0}\subset U_{1}\subset\cdots$ be an increasing sequence of elements in $\mathcal{U}$, where each $U_{i}$ is has compact closure in $U$. Then $\bigcup_{i\geq 0}U_{i}$ belongs to $\mathcal{U}$.
If $M$ is a PL manifold and $\operatorname{Disj}\left(M\right)\subset\mathcal{U}$, then we can also replace condition (3) by a similar condition in which the $A_{i}$’s are compact subcomplexes of $U$ for some triangulation of $U$ coming from the PL structure, by using a handle decomposition associated to the triangulation. (See [RS72, 6.9].)
We now arrive at the proof of Theorems 1.9 and 1.10.
@@
[…]
$\theta:\operatorname{colim}_{I\in\mathcal{I}^{\prime}}F\left(U_{I}\right)\to F\left(U\right)$